# Batch Bayesian Optimization for High-Dimensional Experimental Design: Simulation and Visualization  [^1] A footnote for the title

### v1

@@ Abstract
[…]
@@
Batch Bayesian Optimization, Gaussian Process Regression, High Dimensions, Noise Effect, Synthetic Data, Batch-picking Methods
[…]
IR,NMR,UV
{tocentry}
Some journals require a graphical entry for the Table of Contents. This should be laid out “print ready” so that the sizing of the text is correct.
Inside the tocentry environment, the font used is Helvetica 8 pt, as required by Journal of the American Chemical Society.
The surrounding frame is 9 cm by 3.5 cm, which is the maximum permitted for Journal of the American Chemical Society graphical table of content entries. The box will not resize if the content is too big: instead it will overflow the edge of the box.
This box and the associated title will always be printed on a separate page at the end of the document.
@@ 1 Introduction
Optimizing from a rough process towards a fine-tuned set of process parameters has been the hallmark of success throughout human history. We can think about the shaping of tools during the stone age, the choice of the copper-tin composition in the bronze age, and the carbon content, furnace temperatures, and quenching conditions in the iron age. As our engineering capabilities have advanced throughout the industrial and the modern age, the number of parameters that can be changed during engineering processes have increased. Accelerating the optimization of engineering processes is critical to enable researchers and engineers to allocate more time to infer essential physical and chemical insights from experiments, thus facilitating faster scientific and industrial progress.
@@
BO has recently emerged as a leading method for the efficient, sequential optimization of black-box functions that are costly to evaluate, 1, 2, 3, 4, 5 such as experimental work involving varying levels of automation.6 In BO, a surrogate model is constructed based on the cumulative data at each optimization iteration, and the next point(s) in the domain to test are selected based on the surrogate model according to a user-chosen acquisition policy. The choice of acquisition function and its exploration hyperparameter determines the balance between the exploitation of areas that have previously performed well and the exploration of the unknown regions. While the concept of BO is simple, i.e. to find the input that maximizes the black-box function, many subtleties of BO can befuddle new users.
@@
First, BO operation is highly affected by its selected parameters. Gaussian process regression (GPR)7, 8 is the most popular BO surrogate model because the uncertainty of the model is automatically produced along with the posterior mean. The choice of GPR kernel and its hyperparameters, amplitude and lengthscales, represents the user’s beliefs on what the objective function looks like. Additionally, the GPR noise variance setting should reflect the noise level in the experiment. However, GPR is frequently used for materials science optimization problems that are so novel that the prior knowledge on the properties of the underlying objective function – e.g., its range, the domain landscape, noise level, or the number of optima – are not known beforehand.
@@
The choice of acquisition function also presents another confusion to new users. Various acquisition functions have been presented in literature. Most BO literature, especially papers that are more mathematical and theoretical, use expected improvement (EI).9, 10 In contrast, the upper confidence bound (UCB) 11, 12 is simple and intuitive and is shown to be effective in experimental problems.13, 14 Users often do not know which one to use or how to set the hyperparameter that controls the tradeoff between exploitation and exploration.
@@
Second, material science optimization often depends on many materials components and processing conditions, i.e., they are high-dimensional problems. Since the input parameters may not be independent of each other, it is pertinent to optimize them simultaneously to understand the interactions and dependencies better. With the increasing adaptation of automated/autonomous experiments, the correct implementation of high-dimensional BO algorithms becomes especially critical. However, up until recently, most BO implementations in materials science have only been discussed in low-dimensional, specifically, two- to three-dimensional BO. To our knowledge, the highest-dimensional published BO in experimental materials science work is in five dimensions 15, 16, and there are only a few works in four dimensions 17, 18, 19.
Due to the high dimensionality, the optimization progression is challenging to visualize; hence, evaluating the algorithm-guided experiment loops’ success (or failure) is not easy. A widely used approach in the literature is to use the model’s optimum value for the objective function as the optimization metric. However, this single metric for the BO is not capable of describing the results of the optimization campaign in detail. Furthermore, this presumes that the inputs that produce the optimized objective value of the model are the correct ones to produce the ground truth (GT) optimum, an assumption not yet verified.
Two large differences exist between experimental and simulation work: (1) experiments are often performed in batches, i.e., processing multiple samples at once, to save materials cost or time, and (2) experimental data contain aleatoric uncertainties manifesting as noise. Most BO work are performed for sequential optimization, i.e., picking one sampling condition for the next experiment based on the highest acquisition function value. When performing batch BO, after selecting the first next point at the highest acquisition function value, various selection strategies, categorized as penalizing, exploratory, or stochastic,20, 21, 22, 23 have been proposed to address how to select the remaining points so that they are meaningfully different from the first point and each other. This is challenging because we do not gain any new information between selecting the first and the rest of the batch. Noise is inevitable in experiments and may have a substantial impact to the outcomes in high-dimensional optimization tasks.24, 25, 26, 27, 28 Unfortunately, most of the popular BO algorithm choices implemented in open BO code repositories have been developed and benchmarked in scenarios with negligible experimental noise. Therefore, it is unclear if the same choices of algorithm, acquisition function, or hyperparameter perform well in noisy, high-dimensional experimental tasks encountered in materials science. A systematic evaluation of the effects of noise on optimization outcomes is needed. In simulated BO studies, noise is commonly added based on the proportion of the GT objective value (y) maximum. Depending on the shape of the objective function, this might not reflect the general signal-to-noise level in experiments.
Finally, before implementing BO in an experimental campaign, it is good practice to test BO in a simulation environment using synthetic data, for pedagogical and troubleshooting purposes. In this work, we develop a framework to visualize BO step by step, as an evaluation tool for simulation environments and as a debugging tool for experiments. We showcase an example of simulated data with increasing noise, evaluating optimization strategies as a function of noise magnitude. In our demonstration, we implement batch BO using the Emukit package to find the optimum to two different types of 6-dimensional (6D) objective functions: a needle-in-a-haystack function and a function with a local optimum nearly degenerate with the global optimum. 29, 30
@@ 2 Method
2.1 Synthetic objective functions for benchmarking and analysis of Bayesian optimization
[…]
font=footnotesize Refer to caption
Figure 1: : Visualization of the 3D representation of the (a) Ackley function ground truth, where x1,x2 variables are projected in 3D representation, and (b) Hartmann function ground truth, where x1,x3 variables are projected in 3D representation. The global maximum is labeled as the ‘GT max’ (cyan cross for Ackley, black cross for Hartmann) both functions and Hartmann function 2nd maxima labeled as the ’2nd True Max’ (red star).
We consider two synthetic test functions: (1) an Ackley function31
@@ 2.1 Synthetic objective functions for benchmarking and analysis of Bayesian optimization
where the dimension $d$ = 6, the range of $A({\bf X})$ is $[-22.3,0]$ whereas the domain ${\bf X}$ is a 6D hypercube spanning $[-32.768,32.768]$ along each dimension. Note that compared to Ref. 31, we changed the sign of the Ackley function so that the optimization problem is a maximization problem. The Ackley function, illustrated in Fig.1(a)(only x1,x2 pair variables projected on Ackley function GT) and Fig. S1(all pairs of variables projected on Ackley function GT), has one sharp global maximum $A({\bf X}_{\rm max})=A({\bf 0})=0$ at the origin. The Ackley function thus represents a needle-in-the-haystack-type problem. Functions of this type are also referred to as heterogeneous functions,32, 33 where the convex region around the maximum occupies a minuscule fraction of the parameter space. Specifically, greater than 99.99% of the hypervolume of the domain has $f({\bf X})<-15$, or, equivalently and more precisely, when randomly sampling ${\bf X}$, the probability of finding $f({\bf X})>-15$ is less than $4\times 10^{-5}$.
(2) a Hartmann function31
@@
where $d$ = 6, $\alpha=(1.0,1.2,3.0,3.2)^{T}$ , $A_{ij}$ and $P_{ij}$ are $4\times 6$ matrices defined in Supplementary section 1. The range of $H({\bf X})$ is $[0,3.32237]$ whereas the domain ${\bf X}$ is a 6D unit hypercube spanning from $[0,1]$ along each dimension. We visualize the Hartmann function in Fig.1(b)(only x1,x3 variables pair projected on Hartamnn function GT) and Fig. S2 ( all pairs of variables projected on Hartmann function GT). The Hartmann function has one global maximum $H({\bf X}_{\max})=3.32237$, located at ${\bf X}_{\max}=(0.20169,0.150011,0.476874,$ 0.275332,0.311652,0.6573)0.275332,0.311652,0.6573)0.275332 , 0.311652 , 0.6573 ). A second local maximum $H({\bf X}_{\max,2})=3.20452$ is found at ${\bf X}_{\max,2}=(0.20169,0.150011,0.476874,$ 0.275332,0.311652,0.6573)0.275332,0.311652,0.6573)0.275332 , 0.311652 , 0.6573 ). The gradients near the maximum of the Hartmann function change much more gradually compared to the Ackley function, and we say that the Hartmann function is non-heterogeneous. The Hartmann function represents a shallow maximum and a landscape where optimization can easily get trapped in a local maximum far away from the global maximum.
2.2 Main parameters of Bayesian optimization considered during benchmark and analysis
@@
The first step in building a GPR model in BO is choosing a utility function $u({\bf D})$, where a higher $u$ indicates that the dataset ${\bf D}$ is of higher quality and better able to specify where the global maximum is.34, 35 The most naive choice for the utility function is $u({\bf D})={\rm Max}(y)$, however, this utility function can improve in the presence of noise which is not meaningful. Instead, we use the utility function
@@
where, $\mu_{\bf D}$ is the predicted mean computed using ${\bf D}$, and the maximum (${\rm Max}$) is considered over all ${\bf X}$ in ${\bf D}$. Eq. 3 approaches $u({\bf D})={\rm Max}(y)$ in the absence of noise, but will be a better utility function in the presence of noise.
@@
$\rm UCB({\bf X}|\mu_{\bf D},(\sigma_{\bf D})^{2})=\rm\mu_{\bf D}({\bf X})+\beta(\sigma_{\bf D}({\bf X}))^{2},\beta>0$ (6)
where $\beta$ represents an exploration hyperparameter. We selected UCB and EI because they are used widely, have been implemented in multiple BO code packages, and can tune exploration and exploitation easily via hyperparameters $\beta$ and $\xi$, respectively. In this work, we benchmark both acquisition functions with a range of exploration/exploitation hyperparameter values to evaluate how they affect BO learning outcome.
To mimic experiments in fabricating/testing of multiple samples in a round, we perform batch BO with a batch size of four. We compare three batch-picking methods: Local Penalization (LP),36 Kriging Believer (KB),37 and Constant Liar (CL),38, 39 representing penalizing, exploratory, stochastic strategies and commonly implemented in open BO packages. 40 Although BO was created as a sequential optimization tool, employing batch evaluation is critical to leverage parallelization.

### v3

@@ Abstract
Bayesian Optimization (BO) is increasingly used to guide optimization tasks in materials science research. To emulate real-world experiments, we perform batch BO simulation of two test functions with six design variables and varying noise levels. A needle-in-a-haystack landscape (Ackley function) characterizes searching for a desired but unusual materials property, and a landscape with nearly degenerate optima (Hartmann function) is typically of materials fabrication process. We test different acquisition functions, batch picking methods, and exploration hyperparameter values, and present learning curves for various performance metrics and visualizations to effectively track the optimization progress. Optimization outcomes are shown to depend on noise and the search landscape. Conducting synthetic studies enables researchers to estimate experimental budgets before transitioning to the inherent uncertainties of real experiments. The results and methodology of this study facilitate a greater utilization of BO in guiding experimental materials research, specifically in optimization problems with multiple design variables.
@@
Batch Bayesian Optimization; Materials and Process Optimization; Materials Science Experiments; Synthetic Data; Noise Effect; Problem Landscape
@@ 1 Introduction
Throughout history, breakthroughs in materials science and engineering have relied on the optimization of a rough fabrication process toward a fine-tuned set of process parameters. We can think about the shaping of tools during the stone age, the choice of the copper-tin composition in the bronze age, and the carbon content, furnace temperatures, and quenching conditions in the iron age. As our engineering capabilities advanced throughout the industrial and modern era, the number of design parameters that can be changed during material processing has increased. Identifying optimal values in the input variables will enable the discovery and production of new materials with the desired properties to be faster. Thus, accelerated optimization will enable rapid advancements in materials research.
Machine learning provides powerful new tools for optimization tasks. Bayesian optimization (BO) [1, 2, 3] has recently emerged as the leading method for efficient sequential optimization of black-box functions that are costly to evaluate, situations often encountered in materials science research. [4, 5, 6] Experiments related to materials development — both manual ones and those including varying levels of automation — are typical examples of such costly optimization tasks.
Noise is inevitable in experiments and may have a substantial impact to the outcomes in high-dimensional optimization tasks.[7, 8, 9, 10, 11] Unfortunately, most of the popular BO algorithm choices implemented in open BO code repositories have been developed and benchmarked in scenarios with negligible noise. Therefore, it is unclear if the same choices of algorithm, acquisition function, or hyperparameter perform well in noisy, high-dimensional experimental tasks encountered in materials science or how the results vary for different problems. A systematic evaluation of the effects of noise on the outcome is thus needed and investigated in this work. We also discuss how to simulate noise to reflect the general signal-to-noise level in the experiments, as this is relevant for planning BO campaigns in the laboratory.
In this work, we develop a framework to elucidate every step in Bayesian optimization (BO), enabling its use as an evaluation tool in simulations and as a debugging tool for materials science experiments. We select two six-dimensional (6D) test functions to represent commonly encountered optimization problems in materials science: the Ackley function (a needle-in-a-haystack landscape) and the Hartmann function (which has a local minimum close in value to the global minimum). We examine how increasing noise levels systematically affects optimization outcomes across these two contrasting landscapes. Our benchmarking studies investigate how the choice of optimization metrics, acquisition functions, hyperparameters, and batch-selection methods affect BO outcomes. We also present methodology to visualize the optimization progress in such high-dimensional problems, putting both inputs and objective at equal footing.
Testing BO in simulation with synthetic data before experiments helps researchers learn the method, troubleshoot issues, and verify their experimental budget is adequate. By systematically examining each component of the BO framework, this paper addresses key barriers that materials scientists face when implementing BO. This work bridges the gap between theoretical machine learning literature and practical applications by demonstrating how to build robust, high-dimensional, and noise-aware BO workflows for real experiments, enabling materials scientists to confidently adopt BO as a practical tool for their research.
2 Bayesian Optimization Background
@@
In BO, data are collected iteratively with model guidance. The aim is to maximize the values of an objective function (e.g., mechanical strength of additively manufactured samples [6]) by determining the optimal values for each design variable (such as the angle and diameter of support bars within the structure [6]). There are two key parts in the BO method: a surrogate model and an acquisition function. A surrogate model is rebuilt using the accumulated dataset at each iteration of optimization. The next point(s) to acquire new data are selected from the domain consisting of the design variables based on the surrogate model according to a user-chosen acquisition policy. Specifically, the selection of the next point is performed through evaluating the acquisition function. The choice of acquisition function and its exploration hyperparameter determines the balance between the exploitation of areas that have previously performed well and the exploration of the unknown regions. While the concept of BO is simple, i.e., to find the input values (aka design variables) that maximizes the unknown black-box function, many subtleties of BO can befuddle new users.
It is important to choose a function that represents the search landscape of the research problem. Many materials optimization problems have the needle-in-the-haystack landscape which we emulate here with the Ackley function.[12] Some examples of needle-in-the-haystack landscapes include searching for auxetic materials with a negative Poisson’s ratio,[13] materials with high thermoelectric figure of meric (ZT),[14] and high-entropy alloys that exhibit both high strength and high ductility.[15] The common thread is the search for a highly unusual also highly desired materials properties, which is not common in nature. The Ackley function has its optimal value centered at the origin with a rapid drop-off to fluctuating background values that occupy most of the search space, which represent that the optimal objective value is highly sensitive to small changes in input parameters.
In contrast, false optimal problems are more commonly encountered in process optimization problems,[12] which we emulate using the Hartmann function. In such problems, the optimization space may contain multiple competing high-performing configurations (local optima) alongside a global optimum, posing challenges for efficient search. Examples of this type of landscapes include optimizing deposition parameters for perovskite solar cells,[16] searching for stable perovskite composition,[17], synthesizing silver nanoparticle by microfluidics,[18] and identifying print parameters for enhanced output quality.[19] These types of problems are well represented by the Hartmann function, characterized by gradual variations but containing deceptive local optima that may mislead optimization algorithms.
While many materials optimization problems tend to have a needle-in-a-haystack landscape and many process optimization problems exhibit a false optimal landscape,[12] the division is not strict. For example, in a case of molecular discovery and drug design, Aldeghi showed that the structure-properties relationship landscapes are dramatically different for half-life, hydration free energy, and hepatocyte clearance.[20] In high entropy materials, the properties that follow the rule of mixtures such as bulk modulus and thermal expansion coefficient would likely have a Hartmann-type landscape, while those that depend on specific microstructures, such as high thermal conductivity, would be likely searching for a needle in a haystack.[21] Thus, the choice of which function to adopt as the surrogate model for the experiment will rely on the domain expertise of the researchers. Nevertheless, we believe that Ackley and Hartmann functions cover most materials science problems. Therefore, we systematically investigate how BO performs within Ackley and Hartmann landscape types.
@@
BO operation is highly affected by its selected hyperparameters. Gaussian process regression (GPR) [22, 23] machine learning model is the most popular BO surrogate model because it is a probabilistic model that outputs both the predicted value of the objective function (called in this context as posterior mean) and the uncertainty of the prediction (called in this context as posterior uncertainty). Thus, it is straightforward to exploit areas where predicted values are promising and explore areas where uncertainty is high. The choice of the GPR kernel and its hyperparameters — amplitude and lengthscales — represents the user’s beliefs on what the objective function looks like. Additionally, the GPR noise variance setting should reflect the noise level in the experiment. However, GPR is frequently used for materials science optimization problems so novel that the properties of the underlying objective function — e.g., its range and general shape, the number of optima, or the noise level — are not known beforehand.
@@
The choice of acquisition function presents another confusion to new users due to wide selection in literature and BO code packages. Expected improvement (EI)[24, 25] and upper confidence bound (UCB) [26, 27] are two commonly used ones. Most BO literature — especially those that are mathematical and theoretical — uses EI. In contrast, UCB has been shown to be effective in experimental problems.[28, 29] Users are challenged also by how to set the exploration hyperparameter value, which controls the tradeoff between exploitation and exploration in the acquisition function.
@@
Additionally, practical material science optimization tasks often depend on many materials components and processing conditions, i.e., they are multi-variable problems, which is called high-dimensional problems in machine learning. Since the input parameters may not be independent of each other, it is pertinent to optimize them simultaneously to understand the interactions and dependencies better.[30] With the increasing adaptation of automated/autonomous experiments, the correct implementation of high-dimensional BO algorithms becomes especially critical because human supervision is fragmentary in these cases.
Due to experimental complexity and sampling limitations, most non-automated experiments, still the most widely used method, involve three to six input parameters. This is reflected in the experimental BO literature to date which has been limited to low-dimensional problems. For instance, Gongora et al. applied BO in a 4D space to optimize the toughness of mechanical structures by varying the number of columns, the radius and wall thickness of a column, and the twist angle.[6] Xu et al. optimized flexible solar cell efficiency by tuning 2 precursor variables (concentration and additive volume) and 2 processing variables (photonic curing intensity and length the light is on).[29] Similarly Xie et al. tuned two materials variables (molar ratio and volume) and two process variables (DC voltage and reaction duration) for metal–organic framework synthesis.[31] Deneault et al. used BO to optimize four syringe extrusion parameters (prime delay, print speed, x- and y-position) for additive manufacturing.[19] Mekki-Berrada et al. explored five microfluidic parameters for silver nanoparticle synthesis.[18] In this study, we systematically investigate BO performance in six dimensions (6D) using Ackley and Hartmann test functions as benchmarking for materials science optimization problems.
While 6D is already beyond many experimental BO reports, the number of input variables may span to 8–15 in autonomous labs. Scaling to such regimes introduces new challenges, including kernel decay and degraded uncertainty quantification in GPR models. However, dimensionality reduction through feature selection or extraction can lower the number of important features to a practical level. These can be implemented using ARD kernels [22] to suppress irrelevant inputs or scalable surrogates such as random forests[32]. These provide a pathway to adapt our framework for more complex, higher-dimensional materials systems.
Two large differences exist between experimental and synthetic data optimizations that are relevant for designing experimental BO: (1) experiments are often performed in batches, i.e., processing multiple samples at once, to save materials cost or time; and (2) real-world experimental data contain noise. Most BO work in general machine learning literature are performed for sequential optimization, i.e., picking a single sampling point for the next experiment based on the highest acquisition function value at each iteration of BO. When performing batch BO, the remaining points after the first pick should be selected to be meaningfully different from the already-sampled points; this is not straightforward because we do not gain any new information between selecting the first point and the rest of the batch. Generally there are two approaches for batch picking: serial and parallel.[33] In the serial approach, the first pick is performed by choosing the maximal acquisition function value, the same as non-batch BO, and selecting a strategy, categorized as penalizing, exploratory, or stochastic,[34, 35, 36, 37] to pick the rest of the batch. More advanced true parallel algorithms integrate over a joint probability density function using a generalized non-batch acquisition function.[38] In this work, we focus on the serial approach since it is more intuitive and straightforward and does not require advanced machine learning knowledge. We examine three batch-picking methods: Local Penalization (LP),[34] Kriging Believer (KB),[35] and Constant Liar (CL),[36, 37] representing penalizing, exploratory, stochastic strategy, respectively, and commonly implemented in open BO packages.[39] Yet, their performance across different landscape, acquisition function, and noise level is not clear.
There are many kinds of noise in real-world problems, but in general, they can be characterized as random fluctuations and time-dependent variance. The former affects the precision of the measurements and is the predominant variation for samples made in one single batch, i.e., at the same time. The latter is often seen in samples made in different batches, which may be made at different times, by different researchers, or caused by tool degradation, and determines the accuracy of the result. In this work, we will only focus on random noise because they are present in all experimental data.
Because of the high dimensionality, the optimization progress is challenging to visualize; hence, assessing the progress of the algorithm-guided experiments is not straightforward. A widely used approach in the literature is to monitor the surrogate model’s optimum objective value . However, this single metric for the BO is not capable of describing the results of the optimization campaign in detail. Furthermore, this approach focuses on the output ($y$), while in experiments, the researchers can only controll the design variable input values ($\bf X$}). Implicitly there is an assumption that convergence of the output implies convergence of the input, an assumption not verified until the suggested point has actually been characterized. In this work, we track the optimization progress for both $\bf X$} and $y$ and calculate their instantaneous and cumulative regrets.
@@ 3 Method
3.1 Synthetic Objective Functions for Benchmarking and Analysis of Batch Bayesian Optimization
Figure 1: : Visualization of the 3D representation of the ground truth of (a) Ackley function, where x1,x2 variables are projected in 3D representation, and of (b) Hartmann function, where x1,x3 variables are projected in 3D representation. All the other input variables are chosen so that the maximum A(X) for a given (x1, x2) is shown for Ackley and the maximum H(X) for a given (x1, x3) is shown for Hartmann. The global maximum is labeled as the ‘GT max’ (cyan cross for Ackley, black cross for Hartmann) both functions and Hartmann function 2nd maxima labeled as the ’2nd True Max’ (red star).
To implement test scenarios relevant in materials science, we consider two different types of synthetic test functions: (1) an Ackley function[40]
@@ 3.1 Synthetic Objective Functions for Benchmarking and Analysis of Batch Bayesian Optimization
where the dimension $d$ = 6, the range of $A({\bf X})$ is [22.3,0][-22.3,0] whereas the domain $\bf X$} is a 6D hypercube spanning [32.768,32.768][-32.768,32.768] along each dimension. Note that compared to Ref. [40], we changed the sign of the Ackley function so that the optimization problem is a maximization problem. The Ackley function, illustrated in Fig.1(a)(only x1,x2 pair variables projected on Ackley function GT) and Fig. S1(all pairs of variables projected on Ackley function GT), has one sharp global maximum $A({\bf X}_{\rm max})=A({\bf 0})=0$ at the origin. The Ackley function thus represents a needle-in-the-haystack-type problem. Ackley-type functions are sometimes called heterogeneous functions,[41, 42] where the convex region around the maximum occupies a minuscule fraction of the parameter space. Specifically, greater than 99.99% of the hypervolume of the domain has $f({\bf X})<-15$, or, equivalently and more precisely, when randomly sampling $\bf X$}, the probability of finding $f({\bf X})>-15$ is less than $4\times 10^{-5}$. Some examples of such problem landscapes include searching for auxetic materials with a negative Poisson’s ratio,[13] materials with high thermoelectric figure of meric (ZT),[14] and high-entropy alloys that exhibit both high strength and high ductility.[15] The common thread is the search for a highly unusual also highly desired materials properties, which is not common in nature.
(2) a Hartmann function[40]
@@
where $d$ = 6, $\alpha=(1.0,1.2,3.0,3.2)^{T}$ , $A_{ij}$ and $P_{ij}$ are $4\times 6$ matrices defined in Supplementary section 1. The range of $H({\bf X})$ is $[0,3.32237]$ whereas the domain $\bf X$} is a 6D unit hypercube spanning from $[0,1]$ along each dimension. We visualize the Hartmann function in Fig.1(b)(only x1,x3 variables pair projected on Hartamnn function GT) and Fig. S2 ( all pairs of variables projected on Hartmann function GT). The Hartmann function has one global maximum $H({\bf X}_{\max})=3.32237$, located at ${\bf X}_{\max}=(0.20169,0.150011,0.476874,$ $0.275332,0.311652,0.6573)$. A second local maximum $H({\bf X}_{\max,2})=3.20452$ is found at ${\bf X}_{\max,2}=(0.20169,0.150011,0.476874,$ $0.275332,0.311652,0.6573)$. The gradients near the maximum of the Hartmann function change much more gradually compared to the Ackley function, and we say that the Hartmann function is non-heterogeneous. The Hartmann function represents a shallow maximum and a landscape where optimization can easily get trapped in a local maximum far away from the global maximum. Example optimization problems of this type of search landscape include optimizing deposition parameters for perovskite solar cells,[16] searching for stable perovskite composition,[17], synthesizing silver nanoparticle by microfluidics,[18] and identifying print parameters for enhanced output quality.[19] ]
3.2 Main Parameters of Bayesian Optimization to Be Considered
@@
The utility function $u({\bf D})$ quantifies the quality of the the dataset $\bf D$} where a higher $u$ indicates that $\bf D$} is of higher quality and better able to specify where the global maximum is.[43, 44] The most naive choice for the utility function is the best encountered data point ${\rm Max}(y)$; however, this utility function can yield high values due to noise which is not meaningful. Instead, we use the utility function
@@
where, $\mu_{\bf D}$ is the mean posterior computed using $\bf D$}, and the maximum ($\rm Max$}) is considered over all $\bf X$} in $\bf D$}. Eq. 3 approaches $u({\bf D})={\rm Max}(y)$ in the absence of noise, but will be a better utility function in the presence of noise.
@@
$\rm UCB({\bf X}|\mu_{\bf D},\sigma_{\bf D})=\rm\mu_{\bf D}({\bf X})+\beta\times\sigma_{\bf D}({\bf X}),\beta>0$ (6)
where $\beta$ represents an exploration hyperparameter. In some literature and software packages, the second term of UCB is expressed as $\sqrt{\beta}\sigma$. Here we use the same convention as in Emukit with linear $\beta$. We selected UCB and EI because they are used widely, have been implemented in multiple BO code packages, and can tune exploration and exploitation easily via hyperparameters $\beta$ and $\xi$, respectively. In this work, we benchmark both acquisition functions with a range of exploration/exploitation hyperparameter values to evaluate how they affect BO learning outcome.
To mimic experiments in fabricating/testing of multiple samples in a round, we perform batch BO with a batch size of four. We compare three serial batch-picking methods: Local Penalization (LP),[34] Kriging Believer (KB),[35] and Constant Liar (CL),[36, 37] representing penalizing, exploratory, stochastic strategies and commonly implemented in open BO packages. [39] Although BO was created as a sequential optimization tool, employing batch evaluation is critical to leverage parallelization. KB and CL are greedy versions of more sophisticated true parallel global optimization.[38]